# Yang-Mills Field in the $\kappa$-space-time

### v1

@@ 5 Yang-Mills Lagrangian in $\kappa$-space-time
$\hat{L}=-\frac{1}{2}(1-\frac{ap_{0}}{\hbar})^{2}F^{a}_{0i}F^{0ia}-\frac{1}{4}(1-2\frac{ap_{0}}{\hbar})^{2}F^{a}_{ij}F^{ija},$ (5.1)
which is invariant under the transformation that leaves ${F}^{a}_{0i}F^{a0i}$ and ${F}^{a}_{ij}F^{aij}$ invariant, i.e., usual SU(N) gauge transformations. We have already derived two of the Yang-Mills equations (see eq.(4.15) )and eq.(4.16)). The remaining two equations, follow as the Euler-Lagrange equation from the above Lagrangian, and they are
$e^{-\frac{2ap_{0}}{\hbar}}(D_{i}F^{0i})^{a}=0$ (5.2)
$(D_{0}F^{i0})^{a}+e^{-\frac{2ap_{0}}{\hbar}}(D_{j}F^{ij})^{a}=0$ (5.3)
In the limit $a\rightarrow 0$, above equations reduces to the covariant commutative result. It is clear that under the gauge transformation of the vector field,
$A^{a}_{\mu}\rightarrow A^{a}_{\mu}+(D_{\mu}\alpha(x))^{a},$ (5.4)
where the covariant derivative is $(D_{\mu}\alpha(x))^{a}=(\delta^{ac}\partial_{\mu}-ef^{abc}A^{b}_{\mu}){\alpha}^{c}(x)$, the Lagrangian given in eq.(5.1) is invariant, showing the invariance of $\kappa$-deformed Yang-Mills field theory under usual SU(N) transformation.
@@ 6 Lorentz force equation in presence of SU(N)gauge fields
$[\hat{x}_{i},\hat{F}_{0}]=\frac{i\hbar e}{m}F_{i0}-\frac{a}{\hbar}\Big{(}{\frac{i\hbar e}{m}}F_{i0}p_{0}-\frac{i\hbar}{m}x_{j}\frac{\partial\partial^{j}F_{0}}{\partial\dot{x}^{i}}p_{0}-\frac{i\hbar}{m}\dot{x}_{j}\frac{\partial(\frac{\partial F_{0}}{\partial\dot{x}_{j}})}{\partial\dot{x}^{i}}p_{0}\Big{)}$ (6.2)
@@
From eq.(6.2) and eq.(6.3), we equate the commutative and non-commutative part separately. Equating commutative part gives,
$\frac{i\hbar e}{m}F_{i0}=-\frac{i\hbar}{m}\frac{\partial F_{0}}{\partial\dot{x}^{i}}$ (6.4)
which on integration gives,
@@
$-\frac{i\hbar}{m}\frac{\partial\delta F_{0}}{\partial\dot{x}^{i}}=-\frac{1}{\hbar}\bigg{[}{\frac{i\hbar e}{m}}F_{i0}p_{0}-\frac{i\hbar}{m}x_{j}\frac{\partial\partial^{j}F_{0}}{\partial\dot{x}^{i}}p_{0}-\frac{i\hbar}{m}\dot{x}_{j}\frac{\partial(\frac{\partial F_{0}}{\partial\dot{x}_{j}})}{\partial\dot{x}^{i}}p_{0}\bigg{]}$ (6.6)
@@
and equating the commutative part of eq.(6.10) and eq.(6.11) gives,
$\frac{i\hbar e}{m}F_{ij}=-\frac{i\hbar}{m}\frac{\partial F_{j}}{\partial\dot{x}^{i}}$ (6.12)
which on integration gives
@@
$-\frac{i\hbar}{m}\frac{\partial\delta\bar{F}_{j}}{\partial\dot{x}^{i}}=\frac{i}{m}\bigg{[}x_{k}\frac{\partial\partial^{k}F_{j}}{\partial\dot{x}^{i}}p_{0}+\frac{\partial F_{j}}{\partial\dot{x}^{i}}p_{0}+\dot{x}_{k}\frac{\partial(\frac{\partial F_{j}}{\partial\dot{x}_{k}})}{\partial\dot{x}^{i}}p_{0}\bigg{]}$ (6.14)
Using the commutative expression for $F_{j}=eF_{ji}\dot{x}^{i}+eF_{j0}\dot{x}^{0}+G_{j}$ [30], we get
$\frac{\partial\delta\bar{F}_{j}}{\partial\dot{x}^{i}}=-\frac{1}{\hbar}(ex_{k}\partial^{k}F_{ji}p_{0}+eF_{ji}p_{0})$ (6.15)
On integrating the above equation gives,
@@ 7 Conclusion
We have constructed Yang-Mills theory in $\kappa$-deformed space-time, by generalising Feynman’s approach. For this, we have first constructed the Wong’s equation in $\kappa$-deformed space-time. We have obtained the force equation for a particle having isospin in presence of non-abelian gauge field. We have also constructed the Lagrangian for $\kappa$-deformed Yang-Mills theory and showed its invariance under SU(N) gauge transformations. Thus we see that the $\kappa$-deformed Yang-Mills theory obtained has same gauge invariance as its commutative counter part. In our study we have considered corrections valid only upto first order in the deformation parameter $a$.
In our study, by considering the anti-symmetry property of commutators between velocities, we have introduced a second-rank, anti-symmetric tensor, which later found to be the Yang-Mills field strength tensor in $\kappa$-deformed space-time. Using the Jacobi identity repeatedly, we have showed that this second rank, anti-symmetric tensor obey homogeneous Yang-Mills field equation. In evaluating the Jacobi identity, we need commutation relation between Lie algebra valued functions and $\kappa$-deformed velocities, which in turn leads to the commutation relation between generators of Lie algebra and the velocities. This is obtained by using Wong’s equation in $\kappa$-space-time, which we constructed using the $\kappa$-deformed Dirac Hamiltonian. By repeated use of Jacobi identity we have derived two of the homogeneous Yang-Mills field equation in $\kappa$-deformed space-time.
@@
We have also showed that the commutators of the $\kappa$-deformed gauge covariant derivative, $([\tilde{D}_{0},\tilde{D}_{i}]\phi)^{a}$ and $([\tilde{D}_{i},\tilde{D}_{j}]\phi)^{a}$ match with the Yang-Mills field strength tensors $\hat{F}_{0i}$ and $\hat{F}_{ij}$, respectively, by explicit calculations. We have constructed the Lagrangian describing Yang-Mills theory in $\kappa$-space-time using the field strength tensor and the two of the Yang-Mills field equation are derived by finding the corresponding Euler-Lagrange equation. We analysed the symmetry of the system and showed that the Lagrangian is invariant under SU(N) transformation. We have verified that in the commutative limit, i.e., $a\rightarrow 0$, we get the commutative results, as expected.

### v4

@@ 5 Yang-Mills Lagrangian in $\kappa$-space-time
$\hat{L}=-\frac{1}{2}(1-\frac{ap_{0}}{\hbar})^{2}F^{a}_{0i}F^{a0i}-\frac{1}{4}(1-2\frac{ap_{0}}{\hbar})^{2}F^{a}_{ij}F^{aij},$ (5.1)
is invariant under the usual SU(N) gauge transformations. We have already derived two of the Yang-Mills equations (see eq.(4.11) and eq.(4.12)). The remaining two equations, follow as the Euler-Lagrange equation from the above Lagrangian, and they are
$(1-\frac{2ap_{0}}{\hbar})(D_{i}F^{0i})^{a}=0,~{}~{}(D_{0}F^{i0})^{a}+(1-\frac{2ap_{0}}{\hbar})(D_{j}F^{ij})^{a}=0$ (5.2)
In the limit $a\rightarrow 0$, above equations reduces to the covariant commutative result. It is clear that under the gauge transformation of the vector field, $A^{a}_{\mu}\rightarrow A^{a}_{\mu}+(D_{\mu}\alpha(x))^{a}$, where the covariant derivative is $(D_{\mu}\alpha(x))^{a}=(\delta^{ac}\partial_{\mu}-ef^{abc}A^{b}_{\mu}){\alpha}^{c}(x)$, the Lagrangian given in eq.(5.1) is invariant, showing the invariance of $\kappa$-deformed Yang-Mills field theory under usual SU(N) transformation.
@@ 6 Lorentz force equation in presence of SU(N)gauge fields
$[\hat{x}_{i},\hat{F}_{0}]=\frac{i\hbar e}{m}F_{i0}-\frac{a}{\hbar}\Big{(}{\frac{i\hbar e}{m}}F_{i0}p_{0}-\frac{i\hbar}{m}x_{j}\frac{\partial(\partial^{j}F_{0})}{\partial\dot{x}^{i}}p_{0}-\frac{i\hbar}{m}\frac{\partial F_{0}}{\partial\dot{x}^{i}}p_{0}-\frac{i\hbar}{m}\dot{x}_{j}\frac{\partial(\frac{\partial F_{0}}{\partial\dot{x}_{j}})}{\partial\dot{x}^{i}}p_{0}\Big{)}$ (6.2)
@@
From eq.(6.2) and eq.(6.3), we equate the commutative and non-commutative part separately. Equating commutative part gives, $\frac{i\hbar e}{m}F_{i0}=-\frac{i\hbar}{m}\frac{\partial F_{0}}{\partial\dot{x}^{i}}$. which on integration gives,
@@
$-\frac{i\hbar}{m}\frac{\partial\delta F_{0}}{\partial\dot{x}^{i}}=-\frac{1}{\hbar}\bigg{[}{\frac{i\hbar e}{m}}F_{i0}p_{0}-\frac{i\hbar}{m}x_{j}\frac{\partial(\partial^{j}F_{0})}{\partial\dot{x}^{i}}p_{0}-\frac{i\hbar}{m}\dot{x}_{j}\frac{\partial(\frac{\partial F_{0}}{\partial\dot{x}_{j}})}{\partial\dot{x}^{i}}p_{0}\bigg{]}$ (6.5)
@@
and equating the commutative part of eq.(6.9) and eq.(6.10) gives, $\frac{i\hbar e}{m}F_{ij}=-\frac{i\hbar}{m}\frac{\partial F_{j}}{\partial\dot{x}^{i}}$. which on integration gives
@@
$-\frac{i\hbar}{m}\frac{\partial\delta\bar{F}_{j}}{\partial\dot{x}^{i}}=\frac{i}{m}\bigg{[}x_{k}\frac{\partial(\partial^{k}F_{j})}{\partial\dot{x}^{i}}p_{0}+\frac{\partial F_{j}}{\partial\dot{x}^{i}}p_{0}+\dot{x}_{k}\frac{\partial(\frac{\partial F_{j}}{\partial\dot{x}_{k}})}{\partial\dot{x}^{i}}p_{0}\bigg{]}$ (6.12)
Using the commutative expression for $F_{j}=eF_{ji}\dot{x}^{i}+eF_{j0}\dot{x}^{0}+G_{j}$ [32], we get, $\frac{\partial\delta\bar{F}_{j}}{\partial\dot{x}^{i}}=-\frac{1}{\hbar}(ex_{k}\partial^{k}F_{ji}p_{0}+eF_{ji}p_{0})$. On integrating the above equation gives,
@@ 7 Conclusion
We have constructed Yang-Mills theory in $\kappa$-deformed space-time, by generalising Feynman’s approach. For this, we have first constructed the Wong’s equation in $\kappa$-deformed space-time. We have obtained the force equation for a particle having isospin in presence of non-abelian gauge field. We have also constructed the Lagrangian for $\kappa$-deformed Yang-Mills theory and showed its invariance under SU(N) gauge transformations. Thus we see that the $\kappa$-deformed Yang-Mills theory obtained has the same gauge invariance as its commutative counter part. In our study, we have considered corrections valid only up to first order in the deformation parameter $a$.
In our study, by considering the anti-symmetry property of commutators between velocities, we have introduced a second-rank, anti-symmetric tensor, which was later found to be the Yang-Mills field strength tensor in $\kappa$-deformed space-time. Using the Jacobi identity repeatedly, we have showed that this second rank, anti-symmetric tensor obey homogeneous Yang-Mills field equation. In evaluating the Jacobi identity, we need commutation relation between Lie algebra valued functions and $\kappa$-deformed velocities, which in turn leads to the commutation relation between generators of Lie algebra and the velocities. This is obtained by using Wong’s equation in $\kappa$-space-time, which we constructed using the $\kappa$-deformed Dirac Hamiltonian. Note that the $\kappa$-deformed Dirac equation is constructed [26] using the Dirac derivatives [23] which are written in terms of derivatives with respect to the commutative space-time coordinates and the deformation parameter. Thus the $\kappa$-Dirac equation in momentum space will depend only on commutative momentum $p_{0}$, $p_{i}$. Therefore, the minimal coupling method used to bring in the gauge interaction allows us to introduce only gauge field living in the commutative space-time. Here if by brute force we use $\hat{A}^{\mu}(\hat{x})$, eq.(4.7) will have $a$-dependent term in the RHS while LHS is independent of any, leading to inconsistency. Thus we use $A(x)$ in implementing minimal coupling in eq.(3.8). Note that in the $\kappa$-space-time, it is the time coordinate that does not commute with space-coordinates while space coordinates commute among themselves. This should be contrasted with the Moyal space-time and thus the no-go theorem obtained in Moyal space-time [37] is not applicable to our study. By repeated use of Jacobi identity, then we have derived two of the homogeneous Yang-Mills field equation in $\kappa$-deformed space-time.
@@
We have also showed that the commutators of the $\kappa$-deformed gauge covariant derivative, $([\tilde{D}_{0},\tilde{D}_{i}]\phi)^{a}$ and $([\tilde{D}_{i},\tilde{D}_{j}]\phi)^{a}$ match with the Yang-Mills field strength tensors $\hat{F}_{0i}$ and $\hat{F}_{ij}$, respectively, by explicit calculations. We have constructed the Lagrangian describing Yang-Mills theory in $\kappa$-space-time using the field strength tensor and two of the Yang-Mills field equation are derived by finding the corresponding Euler-Lagrange equation. We analysed the symmetry of the system and showed that the Lagrangian is invariant under SU(N) transformation. We have verified that in the commutative limit, i.e., $a\rightarrow 0$, we get the commutative results, as expected.
@@ 8 Acknowledgement
Data availability statement
No new data were created or analyzed in this study.
Conflict of interest
The authors declare no Conflict of interest.